# How Much Can We Forget about Data Contamination?

### v3

@@ 4.2 Contamination Can be Completely Forgotten
To study the effect of forgetting, we train a 124M parameter model for 15 epochs. Instead of contaminating uniformly over the course of training like in the previous Section 4.1, we perform the contamination between the first and second Chinchilla. [^3] Note that the model is already fairly trained after the first Chinchilla, meaning that the contamination is not very early during training. This is important because there is evidence that observations are more quickly forgotten if the model has not yet learned representations [39, 11, 37]. This is not the setting we are studying here. Figure 2(a) depicts the development of the difference in cross-entropy loss between contaminated and clean benchmark questions over the course of training. We see a strong peak after 2 Chinchilla, which is expected and shows the effect of contamination. What is interesting to us is the rate at which the cross-entropy loss difference decays as we continue training. After training for 1 additional Chinchilla (2.5B tokens for the 124M parameter model), it has already decayed significantly. However, the difference is still visible in Figure 2(a). Figure 2(b) depicts the corresponding accuracy gaps at this point, and we see that all contamination levels still lead to overfitting. As we continue training, the cross-entropy loss difference between contaminated and holdout questions further narrows. From Figure 2(c), which depicts the accuracy gaps after forgetting for a total of 5 Chinchilla, we see that the effect of contamination is eventually completely forgotten in the sense that there is no longer any accuracy difference between contamination and holdout benchmark questions.
@@
The result that contamination can be completely forgotten is in contrast to some previous work on forgetting which have found that forgetting approaches a stable baseline Tirumala et al. [66, Figure 10], or that certain examples are never forgotten [67]. To understand this difference, observe that many previous works on forgetting have not trained on a continuous stream of data. Instead, they have trained on the same training set for multiple epochs. Consequently, we modify our forgetting experiment to repeatedly train on the same 100M tokens after the second epoch. The result of this experiment is depicted in Figure 2(d) and should be compared to Figure 2(a). Interestingly, this simple modification causes the effect of forgetting to stabilize at a level strictly larger than zero. We conclude that exposure to novel data is important for forgetting, an observation similar to Jagielski et al. [39].
@@
4.3 Contamination and Forgetting in OLMo-7B
In the previous sections, we trained small GPT-3 models from scratch. In this Section, we complement this analysis by continual pre-training from an intermediate OLMo-7B checkpoint [30]. Similar to the analysis in Section 4.2, we insert the benchmark data at a specific point into the training data and then measure the subsequent forgetting. Unlike in the previous Section, we now insert the entire benchmark data – we already have a “clean” baseline from the original OLMo-7B training run. We insert each benchmark question four times and contaminate with four different benchmarks: HellaSwag [72], WinoGrande 62, ARC-Easy [19], and PiQA [7].
Figure 3 depicts the result of the experiment. The leftmost point in every plot corresponds to the uncontaminated model, and the following four points depict the effect of contamination. The plots then depict how the effect of contamination is forgotten as we continue training. Similar to the results in the previous sections, the immediate effect of contamination is significant, leading to an average accuracy increase of 17 percentage points across the different benchmarks. At the same time, the effect of contamination decays considerably as we continue training. After continual pre-training for 13% of the remaining training time, the evaluation on WinoGrande and ARC-Easy is no longer significantly different from the uncontaminated evaluation. On HellaSwag and PiQA, the effect of contamination has decayed to approximately 2 percentage points of overfitting.
Supplement Figure 7 depicts the same experiment with OMLo-1B. The behaviour of the 1B parameter model is qualitatively similar to the 7B parameter model seen in this section. However, forgetting for the 1B parameter model occurs faster, corroborating the results from Section 4.1.
5 Weight Decay and Forgetting
@@
In the previous Section 4, we have seen that forgetting is an important empirical property of LLM training. Which factors contribute to the phenomenon? In this section, we show that the weight decay parameter and learning rate schedule of the AdamW optimizer play a part in forgetting past training examples. This offers a novel perspective on the weight decay parameter, usually seen in terms of generalization and training stability [69, 74, 44, 3].
@@ 5.1 The Cumulative Weight Decay
Here, $\theta_{t}$ are the model parameters, $\gamma_{t}$ is the learning rate, $\lambda$ is the weight decay parameter, and $\hat{m_{t}}$ and $\hat{v_{t}}$ are first- and second-order moment estimates of the gradient [49, 61]. Denoting the model weights at initialization by $\theta_{0}$, and the adaptive gradient by $\hat{g}_{t}=\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)$, we can iterate (1) and (2) to obtain
@@
Equation (3) shows that the model weights after $t$ gradient steps are a weighted average of the initial model weights and all the adaptive gradient updates up to time step $t$. The weights $w_{t_{1}}^{t_{2}}$ determine the influence of the gradient update at time $t_{1}$ on the model weights at time $t_{2}$. With weight decay, $w_{t_{1}}^{t_{2}}$ decays to zero, as the following proposition describes.
@@
We present a proof in Supplement A.5. What is the relationship between cumulative weight decay and forgetting? The adaptive gradient $\hat{g}_{t_{1}}$ contains the relevant information about the training examples seen at time $t_{1}$. If $\hat{g}_{t_{1}}$ is no longer part of the model weights because $w^{t_{2}}_{t_{1}}$ has decayed to zero, then the model should no longer depend on the examples seen at time $t_{1}$. In other words, the training examples seen at time $t_{1}$ should be forgotten. However, the data seen at time $t_{1}$ still influences $\theta_{T}$ via the trajectory of gradient descent, even if the direct contribution of $\hat{g}_{t_{1}}$ has decayed to zero. Because of this, we now investigate the relationship between forgetting and weight decay experimentally.
@@ 5.2 Empirical Forgetting Occurs Faster than the Cumulative Weight Decay
We now study the causal effect of weight decay on forgetting by intervening on the weight decay parameter, keeping everything else constant. We consider the contaminated model from Section 4.2 after two epochs and continue training with four different choices of the weight decay parameter. Figure 4 depicts the result of this experiment. Every plot in Figure 4 depicts two curves. The cumulative weight decay is depicted as a solid line. The empirically observed forgetting for 144 times repeated contamination is depicted as a dashed line (compare Figure 2(a)). From Figure 4, we see that increasing the weight decay parameter increases the empirical rate of forgetting. This is evident from the very different x-axis scales in the different plots. Moreover, the dashed line always lies below the solid line, meaning that empirical forgetting occurs faster than the cumulative weight decay. This is despite the fact that weight decay is not necessary for forgetting (see Supplement Figure 8).
(a) 124M 15x Chinchilla
(b) OLMo 7B
(c) Llama 3 405B
Figure 5: Cumulative weight decay and approximate model weight composition for different large-scale training runs. Top Row: The cumulative weight decay $w_{t_{1}}^{t_{2}}$ as defined in equation (4) for different training runs. The figures depict the decay of the gradient updates for every decile of the training run, indicated in different colors. Bottom Row: The approximate composition of the final model weights in terms of the gradient updates from different deciles of the training run. The deciles are indicated in colors depicted in the legend below the plot.
@@ 5.3 Cumulative Weight Decay in Large-Scale Training
Given the discussion in Section 5.1 and the experimental results in Section 5.2, we hypothesize that example forgetting occurs at least as fast as the cumulative weight decay. Based on this hypothesis, we now gauge the degree of forgetting in large-scale training runs (via the cumulative weight decay). The cumulative weight decay only depends on the optimization hyperparameters and can be cheaply computed. [^4] Ideally, we would want experimental evidence on forgetting in large-scale training, akin to the analysis in Section 4.1. Because this exceeds our experimental budget, we rely on the assumption that forgetting in large-scale training occurs at least as fast as the cumulative weight decay, similar to what we have observed for the smaller model in Section 5.2.
Figure 5 depicts the cumulative weight decay $w_{t_{1}}^{t_{2}}$ for three different models: The 124M parameter model from Section 4.2, OLMo 7B [30], and Llama 3 405B [22] (where we assume the model trained with a weight decay of 0.1). The figure depicts the evolution of the term $w_{t_{1}}^{t_{2}}$ for ten fixed choices of $t_{1}$, one at each decile of the training run. Under the assumption that example forgetting occurs at least as fast as the cumulative weight decay, we can interpret the curves in Figure 5 as forgetting curves that indicate to what degree the data seen at different stages is forgotten as we continue training. For the 124M model depicted in Figure 5(a), even the initialization is not completely decayed towards the end (the blue curve is still significantly larger than zero). For OLMo-7B, the gradient weights of the first 40% of the training data decay to zero until the end of training, meaning that this data is likely forgotten (Figure 5(b)). Llama 3 405B also experiences significant decay of the early gradients (Figure 5(c)).
By integrating $w_{t_{1}}^{T}$, the cumulative weight decay allows us to approximate the composition of the final model weights in terms of the gradient updates from different deciles of training. This is depicted in the bottom row of Figure 5.
@@ 6 Discussion
In this work, we have shown that the impact of data contamination depends on the joint scaling of model, data and contamination — an aspect that has largely been overlooked in the literature [71, 57, 40]. Moreover, the impact of contamination can be forgotten as the scale of the data increases, and this is likely the case in many realistic training runs [30, 25].
We have studied contamination with a focus on the leakage of benchmark questions into the training data. This means our work might be more informative than other works that study contamination in different contexts. At the same time, one has to be careful when extrapolating our results, especially to a privacy setup [12, 39]. This is because empirical forgetting might behave differently for random strings or otherwise uniquely identifiable information [13].
The “contamination problem” as studied in this paper has interesting connections to many areas of machine learning, including privacy [29, 39], data attribution [41], and generalization [9, 33, 53]. The connection to data attribution is especially relevant because we show that the presence or absence of a datapoint in the training data can be irrelevant for the model behavior on that same datapoint.
@@ Acknowledgments
We would like to thank Jonas Geiping and Dirk Groeneveld for helpful discussions about data contamination. This project is supported by the Tübingen AI Center and the German Research Foundation through the Cluster of Excellence “Machine Learning - New Perspectives for Science" (EXC 2064/1 number 390727645).
@@ Appendix A Appendix
A.1 Additional Discussion of Data Contamination Assumptions and Setting

### v4

@@ 4.2 Contamination Can be Completely Forgotten
Figure 3: Contamination and Forgetting in OLMo-1B and OLMo-7B. We contaminate intermediate OLMo checkpoints four times with different benchmarks. This causes an average accuracy increase of 17 percentage points for the 7B model. We then continue pre-training for 13% of the remaining training time (or 25000 gradient steps), leading to a reduction of 87% of the accuracy increase due to contamination. The solid line depicts the result for OLMo-7B, and the dashed line depicts the results for OLMo-1B. To account for the difference in the number of model parameters, we scaled the curve of the 1B parameter model by a factor of 5.9 (the actual parameter ratio between the two models), after which the forgetting curves of both models align remarkably well (the plot depicts approximately 3500 gradient steps for the 1B model, and 25000 gradient steps for the 7B model). Different colors correspond to different benchmarks, and the grey line depicts the clean accuracy before contamination. Supplement Figures 9 and 10 depict the results for both models separately.
@@
To study the effect of forgetting, we train a 124M parameter model at 15x Chinchilla. Instead of contaminating uniformly over the course of training like in the previous Section 4.1, we perform the contamination between the first and second Chinchilla. [^4] Note that the model is already fairly trained after the first Chinchilla, meaning that the contamination is not very early during training. This is important because there is evidence that observations are more quickly forgotten if the model has not yet learned representations (Jagielski et al., 2023; Cao et al., 2024; Huang et al., 2024b). This is not the setting we are studying here. Figure 2(a) depicts the development of the difference in cross-entropy loss between contaminated and clean benchmark questions over the course of training. We see a strong peak after 2 Chinchilla, which is expected and shows the effect of contamination. What is interesting to us is the rate at which the cross-entropy loss difference decays as we continue training. After training for 1 additional Chinchilla (2.5B tokens for the 124M parameter model), it has already decayed significantly. However, the difference is still visible in Figure 2(a). Figure 2(b) depicts the corresponding accuracy differences at this point, and we see that all contamination levels still lead to overfitting. As we continue training, the cross-entropy loss difference between contaminated and holdout questions further narrows. From Figure 2(c), which depicts the accuracy differences after forgetting for a total of 5 Chinchilla, we see that the effect of contamination is eventually completely forgotten in the sense that there is no longer any accuracy difference between contamination and holdout benchmark questions.
@@
The result that contamination can be completely forgotten is in contrast to some previous work on forgetting which have found that forgetting approaches a stable baseline Tirumala et al. (2022, Figure 10), or that certain examples are never forgotten (Toneva et al., 2019). To understand this difference, observe that many previous works on forgetting have not trained on a continuous stream of data. Instead, they have trained on the same training set for multiple epochs. Consequently, we modify our forgetting experiment to repeatedly train on the same 100M tokens after the second epoch. The result of this experiment is depicted in Figure 2(d) and should be compared to Figure 2(a). Interestingly, this simple modification causes the effect of forgetting to stabilize at a level strictly larger than zero. We conclude that exposure to novel data is important for forgetting, an observation similar to Jagielski et al. (2023).
@@
4.3 Contamination and Forgetting in OLMo
In the previous sections, we trained small GPT-3 models from scratch. In this Section, we complement this analysis by continual pre-training from intermediate OLMo-1B and OLMo-7B checkpoints (Groeneveld et al., 2024). Similar to the analysis in Section 4.2, we insert the benchmark data at a specific point into the training data and then measure the subsequent forgetting. Unlike in the previous Section, we now insert the entire benchmark data – we already have a “clean” baseline from the original OLMo training runs. We insert each benchmark question four times, and contaminate with four different benchmarks: HellaSwag (Zellers et al., 2019), WinoGrande Sakaguchi et al., 2021, ARC-Easy (Clark et al., 2018), and PiQA (Bisk et al., 2020).
Figure 3 depicts the result of the experiment. The leftmost point in every plot corresponds to the uncontaminated model, and the following four points depict the effect of contamination. The plots then depict how the effect of contamination is forgotten as we continue training. Similar to the results in the previous sections, the immediate effect of contamination is significant, leading to an average accuracy increase of 17 percentage points across the different benchmarks for the 7B model. At the same time, the effect of contamination decays considerably as we continue training. After continual pre-training for 13% of the remaining training time, the evaluation on WinoGrande and ARC-Easy for the 7B model is no longer significantly different from the uncontaminated evaluation. On HellaSwag and PiQA, the effect of contamination has decayed to approximately 2 percentage points of overfitting. This demonstrates that the effect of forgetting during pre-training is significant, even for a 7B parameter model.
Interestingly, we find that forgetting in OLMo exhibits a scaling behavior. To see this, note that Figure 3 also depicts the result of the forgetting experiment with OLMo-1B. However, we rescaled the forgetting curve of the 1B parameter model by a factor of 5.9, meaning that at any given point on the x-axis, the 7B parameter model has seen 5.9 times as many tokens as the 1B parameter model. [^5] Here we refer to the number of tokens seen in the experiment. The total number of tokens seen by the two models during pre-training up to this point is approximately the same. We choose the factor 5.9 because it corresponds to the actual parameter ratio between the two models (OLMo-1B has 1,176,764,416 parameters, whereas OLMo-7B has 6,888,095,744 parameters). With this rescaling, the forgetting curves of the two models are surprisingly well aligned, suggesting that forgetting with a model that has 5.9 times as many parameters requires 5.9 times as many tokens. Of course, this observation is well aligned with the results on scaling in Section 4.1.
5 Weight Decay and Forgetting
(a) 124M 15x Chinchilla
(b) OLMo 7B
(c) Llama 3 405B
Figure 5: Cumulative weight decay and approximate model weight composition for different large-scale training runs. Top Row: The cumulative weight decay $w_{t_{1}}^{t_{2}}$ as defined in equation (4) for different training runs. The figures depict the decay of the gradient updates for every decile of the training run, indicated in different colors. Bottom Row: The approximate composition of the final model weights in terms of the gradient updates from different deciles of the training run. The deciles are indicated in colors depicted in the legend below the plot. The cumulative weight decay can be easily computed using this Jupyter Notebook.
In the previous Section 4, we have seen that forgetting is an important empirical property of LLM training. Which factors contribute to the phenomenon? In this section, we show that the weight decay parameter and learning rate schedule of the AdamW optimizer play a part in forgetting past training examples. This offers a novel perspective on the weight decay parameter, usually seen in terms of generalization and training stability (Van Laarhoven, 2017; Zhang et al., 2019; Lewkowycz & Gur-Ari, 2020; Andriushchenko et al., 2024).
@@ 5.1 The Cumulative Weight Decay
Here, $\theta_{t}$ are the model parameters, $\gamma_{t}$ is the learning rate, $\lambda$ is the weight decay parameter, and $\hat{m_{t}}$ and $\hat{v_{t}}$ are first- and second-order moment estimates of the gradient (Loshchilov & Hutter, 2019; PyTorch Contributors, 2024). Denoting the model weights at initialization by $\theta_{0}$, and the adaptive gradient by $\hat{g}_{t}=\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)$, we can iterate (1) and (2) to obtain
@@
Equation (3) shows that the model weights after $t$ gradient steps are a weighted average of the initial model weights and all the adaptive gradient updates up to time step $t$. In other words, the weights $w_{t_{1}}^{t_{2}}$ determine the influence of the gradient update at time $t_{1}$ on the model weights at time $t_{2}$. As recently observed by Wang & Aitchison (2025), the weights $\theta_{T}$ can also be understood as an exponential moving average of past gradient updates. With weight decay, $w_{t_{1}}^{t_{2}}$ decays to zero, as the following proposition describes.
@@
We present a proof in Supplement A.8.
What is the relationship between cumulative weight decay and forgetting? The adaptive gradient $\hat{g}_{t_{1}}$ contains the relevant information about the training examples seen at time $t_{1}$. If $\hat{g}_{t_{1}}$ is no longer part of the model weights because $w^{t_{2}}_{t_{1}}$ has decayed to zero, then the model should no longer depend on the examples seen at time $t_{1}$. In other words, the training examples seen at time $t_{1}$ should be forgotten. However, the data seen at time $t_{1}$ still influences $\theta_{T}$ via the trajectory of gradient descent, even if the direct contribution of $\hat{g}_{t_{1}}$ has decayed to zero. Because of this, we now investigate the relationship between forgetting and weight decay experimentally.
@@ 5.2 Empirical Forgetting Occurs Faster than the Cumulative Weight Decay
We now study the causal effect of weight decay on forgetting by intervening on the weight decay parameter, keeping everything else constant. We consider the contaminated model from Section 4.2 after two Chinchilla and continue training with four different choices of the weight decay parameter. Figure 4 depicts the result of this experiment. Every plot in Figure 4 depicts two curves. The cumulative weight decay is depicted as a solid line. The empirically observed forgetting for 144 times repeated contamination is depicted as a dashed line (compare Figure 2(a)). From Figure 4, we see that increasing the weight decay parameter increases the empirical rate of forgetting. This is evident from the very different x-axis scales in the different plots. Moreover, the dashed line always lies below the solid line, meaning that empirical forgetting occurs faster than the cumulative weight decay. This is despite the fact that weight decay is not necessary for forgetting (see Supplement Figure 11).
@@ 5.3 Cumulative Weight Decay in Large-Scale Training
Given Proposition 1 and the experimental results in Section 5.2, we hypothesize that example forgetting occurs at least as fast as the cumulative weight decay. Based on this hypothesis, we now gauge the degree of forgetting in large-scale training runs. The key insight here is that the cumulative weight decay depends only on the learning rate schedule and the weight decay parameter. This means that we can study the cumulative weight decay even in large-scale training runs that far exceed our experimental budget. Ideally, we would of course want experimental evidence on forgetting in large-scale training, akin to the analysis in Section 4.1. Since this is not possible within our experimental budget, we resort to a heuristic analysis of example forgetting via the cumulative weight decay.
Figure 5 depicts the cumulative weight decay $w_{t_{1}}^{t_{2}}$ for three different models: The 124M parameter model from Section 4.2, OLMo 7B (Groeneveld et al., 2024), and Llama 3 405B (Dubey et al., 2024) (where we assume the model trained with a weight decay of 0.1). The figure depicts the evolution of the term $w_{t_{1}}^{t_{2}}$ for ten fixed choices of $t_{1}$, one at each decile of the training run. As discussed above, we can interpret the curves in Figure 5 as heuristic forgetting curves that indicate the degree to which the data seen at different stages is forgotten as we continue training. For the 124M model depicted in Figure 5(a), even the initialization is not completely forgotten at the end of training (the blue curve is still significantly larger than zero). For OLMo-7B, however, the weights of the gradients of the first 40% of training decay to zero until the end of training, meaning that the data seen during this time is likely forgotten (Figure 5(b)). The same holds for Llama 3 405B, where the first 10% of the training data is likely forgotten (Figure 5(c)).
By integrating $w_{t_{1}}^{T}$, the cumulative weight decay even allows us to approximate the composition of the final model weights in terms of the gradient updates from different deciles of training. This is depicted in the bottom row of Figure 5.
@@ 6 Discussion
In this work, we have connected the literature on data contamination (Jiang et al., 2024) with research on forgetting (Tirumala et al., 2022; Jagielski et al., 2023). Through careful controlled experiments, we have shown that the impact of contamination depends on the joint scaling of model, data and contamination — an aspect that has largely been overlooked in the existing literature (Yang et al., 2023; Oren et al., 2024; Jiang et al., 2024). Perhaps surprisingly, we have also seen that the impact of contamination can be completely forgotten, an effect that is especially relevant for training runs that operate in the large-data regime (Groeneveld et al., 2024; Gemma Team, 2024). As a consequence, large-data LLM pre-training also appears to be empirically stable with respect to individual benchmark questions (Bousquet & Elisseeff, 2002). Interestingly, we have also seen a number of ways in which LLM pre-training is different from other learning setups. For example, we have seen that forgetting behaves very differently when training on a continuous stream of novel data as opposed to multi-epoch training on a limited dataset (Tirumala et al., 2022; Muennighoff et al., 2023).
The “contamination problem” as studied in this paper has interesting connections to many areas of machine learning, including privacy (Graves et al., 2021; Jagielski et al., 2023), data attribution (Kirchenbauer et al., 2024), and generalization (Bousquet & Elisseeff, 2002; Hardt et al., 2016; Mania et al., 2019). The connection to data attribution arises from the fact that the presence or absence of a benchmark question in the pre-training data sometimes appears to be irrelevant for the model behavior on that same datapoint, suggesting that it does not make sense to attribute model behavior to individual datapoints in this regime.
An important limitation of our work is that our empirical results on forgetting are restricted to benchmark questions. This means that one has to be careful when extrapolating our results to other contexts, especially to a privacy setup (Carlini et al., 2019; Jagielski et al., 2023). This is because empirical forgetting might behave differently for random strings or otherwise uniquely identifiable information (Carlini et al., 2021).
Impact Statement
This paper studies the empirical properties of training deep neural networks. Because our study concerns the scientific question of model evaluation, we don’t believe it raises significant ethical concerns. That being said, contamination and memorization have broader implications, including copyright and privacy.
@@ Acknowledgments
We would like to thank Jonas Geiping and Dirk Groeneveld for helpful discussion about data contamination. This work is supported by the Tübingen AI Center, the German Research Foundation through the Cluster of Excellence “Machine Learning - New Perspectives for Science” (EXC 2064/1 number 390727645), and the CZS Institute for Artificial Intelligence and Law.
@@ Appendix A Appendix
A.1 Takeaways
Here we list some important takeaways from our work.
  1. 1.
    The impact of data contamination is not fixed, but depends on the scale of model, data, and contamination.
  2. 2.
    If the scale of the data is large compared to the scale of the model and contamination, then the causal effect of contamination can be zero.
  3. 3.
    The interplay of two opposing mechanisms determines the impact of seeing a benchmark question during pre-training. Directly after seeing the question, there is a spike in the likelihood that corresponds to overfitting. Subsequently, there is a forgetting effect, meaning that the impact of the text on the model decays as we continue training.
  4. 4.
    LLM pre-training is fairly stable with respect to individual benchmark questions.
  5. 5.
    LLM pre-training, where the model is continuously exposed to novel data over many gradient steps, differs from multi-epoch training and small-data fine-tuning setups because of forgetting.
  6. 6.
    The weight decay parameter in the AdamW optimizer has a causal effect on forgetting.
  7. 7.
    Cumulative weight decay might be a useful heuristic for forgetting in large-scale training runs.
A.2 Limitations
Here we list some important limitations of our work.
  1. 1.
    The experiments in this paper were conducted only with benchmark questions, and only for the setup of exact contamination.
  2. 2.
    Forgetting might behave differently for random strings or other uniquely identifiable information.
  3. 3.
    Further experiments would be required to validate our results at a larger scale ($>7$B parameters).
A.3 Additional Discussion of Data Contamination Assumptions and Setting